# Decoherence in Exchange-Coupled Quantum Spin Qubit Systems: Impact of Multiqubit Interactions and Geometric Connectivity

### v1

@@ III.2 Entanglement entropy for different geometric connectivity
Because tree configuration can lead to a large system size, our current study focuses specifically on the results of the “node” elementary unit’s tree configuration for a system size of $L=8$. In addition to extracting the decoherence lifetime, we also calculate the entanglement entropy, a highly relevant physical quantity in theoretical studies. We partition the system into a target system $\alpha$ and an environment $\beta$, and obtain the density matrix of the target system by performing a partial trace operation: $\rho_{\alpha}(t)={\rm Tr}_{\beta}\rho(t)$, where $\rho(t)=|\Phi(t)\rangle\langle\Phi(t)|$ represents the density matrix of the entire system at time $t$. The entanglement entropy is defined as
@@
We investigate the influence of varying the deviation $\sigma$ on the decoherence time for a system with $L=8$ qubits. In the case of the “node” elementary unit, as $\sigma$ increases, the decoherence time gradually decreases for all three connectivity types: ring, chain, and tree. However, the ring geometry shows a faster rate of decrease compared to the chain and tree. Despite the variations, the ring consistently outperforms the chain and tree in terms of decoherence time, indicating that the ring geometry is more stable, in line with our previous conclusions.
Moreover, an important observation from the figures is that the chain and tree geometries exhibit similar decoherence times (Fig. 8). This finding highlights the comparable stability of these two connectivity types. Our results help in strengthening our understanding of how different connectivity patterns impact the overall stability and performance of quantum systems. The geometric configuration plays a crucial role in determining the decoherence behavior, which is valuable for designing robust quantum systems and devising quantum error correction strategies.
Figure 8: The decoherence time $T_{2}^{*}$ as a function of $\sigma$ for elementary units A for ring, chain and tree configuration at system size $L=8$. The interacting strength $J_{0}=100/t_{0}$.
@@
(b) Refer to caption (c) Refer to caption
Figure 9: The entanglement entropy $S(t)$ for elementary unit “node” for (a) ring, (b) chain and (c) tree configuration at system size $L=8$. The trend of the entanglement entropy $S(t)$ is very close to that of the return probability $P(t)$, so both can be used to characterize the decoherence lifetime.
@@ IV Conclusions and discussions
In summary, this paper investigates the influence of connectivity on the decoherence time under quasi-static Heisenberg noise. By examining various configurations of elementary units, including rings, chains, and trees, we uncover interesting and counterintuitive results. Specifically, we find that rings exhibit greater stability compared to chains, contrary to the expectation that more links between qubits leads to decreased stability. We attribute this difference to the role of symmetry, which can protect the quantum state from leakage into other subspaces. Additionally, we observe that the “stick” configuration is more stable than the “triangle” configuration, which is expected due to the disparity in their average connectivity. Notably, the disparity in decoherence time between rings and chains is significant, even though their average connectivity is relatively similar. These findings shed light on the complex interplay between connectivity and stability in quantum systems.
@@ Acknowledgements.
Appendix A Dynamic noise

### v2

@@ III.2 Entanglement entropy for different geometric connectivity
Due to the potential for large system sizes in tree configurations, our current study of the tree configuration specifically focuses on the results of the “node” elementary unit for a system size of $L=8$. In addition to extracting the decoherence lifetime, we also calculate the entanglement entropy, a highly relevant physical quantity in theoretical studies. We partition the system into a target system $\alpha$ and an environment $\beta$, and obtain the density matrix of the target system by performing a partial trace operation: $\rho_{\alpha}(t)={\rm Tr}_{\beta}\rho(t)$, where $\rho(t)=|\Phi(t)\rangle\langle\Phi(t)|$ represents the density matrix of the entire system at time $t$. The entanglement entropy is defined as
@@
We investigate the influence of varying the deviation $\sigma$ on the decoherence time for a system with $L=8$ qubits. In the case of the “node” elementary unit, as $\sigma$ increases, the decoherence time gradually decreases for all three connectivity types: ring, chain, and tree. However, the ring configuration shows a faster rate of decrease compared to the chain and tree. Despite the variations, the ring consistently outperforms the chain and tree in terms of decoherence time, indicating that the ring configuration is more stable, in line with our previous conclusions.
Moreover, an important observation from the figures is that the chain and tree configurations exhibit similar decoherence times (Fig. 8). This finding highlights the comparable stability of these two connectivity types. Our results help in strengthening our understanding of how different connectivity patterns impact the overall stability and performance of quantum systems. The geometric configuration plays a crucial role in determining the decoherence behavior, which is valuable for designing robust quantum systems and devising quantum error correction strategies.
Figure 8: The decoherence time $T_{2}^{*}$ as a function of $\sigma$ for elementary units “node” for ring, chain and tree configuration at system size $L=8$. The interacting strength $J_{0}=100/t_{0}$.
@@
(b) Refer to caption
(c) Refer to caption
Figure 9: The entanglement entropy $S(t)$ (upper curves) and the return probability $P(t)$ (lower curves) for elementary unit “node” for (a) ring, (b) chain and (c) tree configuration at system size $L=8$. The trend of the entanglement entropy $S(t)$ is very close to that of the return probability $P(t)$, so both can be used to characterize the decoherence lifetime.
@@ IV Conclusions and discussions
In summary, this paper investigates the influence of connectivity on the decoherence time under quasi-static Heisenberg noise. By examining various configurations of elementary units, including rings, chains, and trees, we uncover interesting and counterintuitive results. Specifically, we find that rings exhibit greater stability compared to chains, contrary to the expectation that more links between qubits leads to decreased stability. We attribute this difference to the role of symmetry, which can protect the quantum state from leakage into other subspaces.
Additionally, we observe that the “stick” configuration is more stable than the “triangle” configuration, which is expected due to the disparity in their average connectivity. Notably, the disparity in decoherence time between rings and chains is significant, even though their average connectivity is relatively similar. These findings shed light on the complex interplay between connectivity and stability in quantum systems.
@@ Acknowledgements.
Appendix A The effect of single-qubit term
In Sec. II, we mentioned that the results shown in the main text are obtained with single-qubit interaction terms (Zeeman terms such as $E_{z}\sigma_{iz}$) ignored. The main reason is that errors incurred in two-qubit gates are much higher than those for single-qubit gates, as pointed out in Johnstun and Van Huele (2021), and crosstalk plays a key role in the two-qubit decoherence process Throckmorton and Das Sarma (2022). In this section, we conduct calculations in presence of single-qubit terms. In a study Di Bartolomeo et al. (2023) on IBM’s Qiskit, the simulated $T_{2}^{*}$ under single-qubit gate operations is ten times longer than that of two-qubit gate operations. This gives us a rough estimate of $E_{z}\approx 0.1J_{0}$. The calculations with $E_{z}=0.1J_{0}$ is shown in Fig. 10, which is compared to the results with $E_{z}=0$. The results from $E_{z}=0.1J_{0}$ and $E_{z}=0$ are rather close, suggesting that our main conclusion remain valid in presence of a resonable low level of single-qubit error terms.
Figure 10: The decoherence time $T_{2}^{*}$ is plotted as a function of $L$ for systems composed of nodes in rings and chains, both with and without single-qubit terms. We take $\sigma=0.4/t_{0}$, $J_{0}=100/t_{0}$ and $E_{z}=0.1J_{0}$.
Appendix B Dynamic noise
@@
Figure 12: Visualization of (a) chain and (b) ring configurations of the elementary unit “stick”. The vertices represent qubits, with gray lines indicating connections within each unit and red lines showing connections between repeated units. The arrows inside the vertices represent the chosen initial states: an upward arrow signifies a spin-up state, while a downward arrow signifies a spin-down state.
Figure 13: Visualization of (a) chain and (b) ring configurations of elementary unit “triangle”.
@@
Appendix C Visualization of chain and ring configurations studied in this work
In this appendix, we present visualizations of chain and ring configurations formed by elementary units “stick” and “triangle,” with the number of qubits, $L$, ranging from 4 to 10. Fig. 12 illustrates the chain and ring configurations of “stick,” while Fig. 13 depicts the chain and ring configurations of “triangle.”
Appendix D Decoherence of GHZ states
Figure 14: The decoherence time $T_{2}^{*}$ as a function of $L$ for the product state and the general GHZ state in different systems.
As discussed in the main text, comparing different geometric configurations constrains our choice of initial states. We have chosen the product state $|\Psi(0)\rangle=|\uparrow\downarrow\cdots\rangle$ as one candidate. Another potential choice is the GHZ state, defined as $|\Psi(0)\rangle_{G}=\frac{1}{\sqrt{2}}(|\uparrow\uparrow\cdots\rangle+|\downarrow\downarrow\cdots\rangle)$. We have conducted calculations for the GHZ state in both ring and chain configurations based on nodes. The results are presented in Fig. 14. It is evident that the ring configuration with nodes exhibits a longer decoherence time than the chain configuration, supporting our main conclusion. Additionally, we note the absence of dips for the ring case at odd $L$, indicating that the “even-odd” effect does not apply to GHZ states.
Finally, we acknowledge that our choice of initial states is not exhaustive. As near-term quantum devices continue to scale up, the interplay between geometric configuration, initial states, and device properties, including decoherence times, emerges as a significant area of interest that merits further investigation.